# zkCraft: Prompt-Guided LLM as a Zero-Shot Mutation Pattern Oracle for TCCT-Powered ZK Fuzzing

### v1

@@ 1 Introduction
Zero-knowledge circuits provide succinct, verifiable proofs while exposing minimal information about private inputs, and they form the foundation of many privacy and scalability systems. Writing correct circuits is difficult because the witness generator that computes private values must align precisely with the constraint system used for proof verification. When this alignment fails, two broad classes of faults appear. Under-constrained faults arise when the constraint system admits execution traces that no valid witness can produce. Over-constrained faults arise when a valid witness-producing execution is rejected by the constraints. Both fault classes have important security and operational consequences in deployed systems and have driven significant recent research efforts  Pailoor et al. (2023); Wen et al. (2024); Cuéllar Gempeler et al. (2025); Xiao et al. (2025); Ron-Zewi and Weiss (2024); Luick et al. (2024).
Prior work addresses these problems from multiple perspectives. Static and pattern-based detectors reveal common syntactic mistakes but often miss semantic discrepancies that manifest only during witness generation or under specific inputs  Chin et al. (2021); Vo et al. (2025); Stronati et al. (2024). Formal verification and SMT-based analyses can provide strong guarantees but typically do not scale to large circuits, and they struggle with the combined search space of program edits and input assignments  Wei et al. (2025); Stephens et al. (2025); Isabel et al. (2024). Dynamic approaches based on fuzzing and metamorphic testing improve practical bug-finding coverage, yet they frequently depend on many expensive solver calls or unguided mutations that produce numerous low-value candidates  Pailoor et al. (2023); Wen et al. (2024); Xiao et al. (2025). Recent work on verifier-side encodings, interactive oracle proofs, and efficient polynomial commitments offers promising primitives to compress or certify search outcomes  Ron-Zewi and Weiss (2024); Zhang et al. (2024b, 2023). In parallel, advances in model-guided mutation and LLM-driven fuzzing have shown that language models can generate semantically rich test patterns and seed mutations for security testing  Meng et al. (2024); Zhang et al. (2024a); Lu et al. (2025); Zhang et al. (2025); Ashkenazi et al. (2025); Li and Shin (2024).
Despite these developments, important gaps remain. Most algebraic and IOP-style techniques have not been integrated with practical program-level mutation strategies, so solver costs remain a bottleneck for end-to-end bug discovery. Conversely, LLM-guided mutation methods improve suggestion quality but they are typically detached from proof-bearing search and therefore can produce proposals that are expensive to validate or that lead to many false positives  Sun et al. (2024); Li and Shin (2024); Zhang et al. (2024a). Compilers and zkVM toolchains introduce additional complexity related to optimization and backend behavior, which further complicates end-to-end detection and increases the cost of exhaustive solver-based validation  Gassmann et al. (2025); Xu et al. (2025); Zou et al. (2025); Xie et al. (2025).
@@ 2.1 Correctness and verification of ZK pipelines and compilers
The correctness of zero-knowledge compilation and circuit-processing pipelines has attracted growing attention because compiler or pipeline bugs can yield accepted yet invalid proofs. Static and dynamic analyses specifically targeting ZK circuit representations and their compilers have been proposed to reveal semantic and logic flaws early in the toolchain. Prior work presents static detectors over circuit dependence graphs and evaluates them on large Circom codebases, demonstrating that semantic checks can recover subtle vulnerabilities in real projects  Wen et al. (2024). Metamorphic and mutation-driven approaches have been adapted to the ZK compiler setting to detect incorrect compilation behaviors by generating semantically equivalent variants and checking for inconsistent outputs  Xiao et al. (2025); Li et al. (2024). Systematic fuzzing of end-to-end ZK pipelines has also been shown to surface logic bugs that elude conventional testing, illustrating the importance of domain-aware test oracles for pipelines that span DSL, compiler, and proof-generation layers  Hochrainer et al. (2025a); Chaliasos et al. (2025).
@@ 2.2 Fuzzing, metamorphic testing, and oracles for ZK systems
Fuzzing and metamorphic testing have been successfully adapted to find pipeline-level and finalization bugs in ZK deployments, including rollup finalization failures and pipeline logic flaws; these techniques emphasize careful oracle design and input-generation strategies to trigger diverse behaviors  Li et al. (2024); Peng et al. (2025). Tools that combine metamorphic relations with fault-injection or targeted mutation expose both soundness and completeness weaknesses in zkVMs and related infrastructures  Hochrainer et al. (2025b); Li et al. (2024). Work that specifically targets the oracle problem in ZK fuzzing highlights the need for proofs or semantic checks that distinguish genuine counterexamples from benign variations  Chaliasos et al. (2025); Hochrainer et al. (2025a). Complementary studies show that bounded, structured mutation templates and algebraic checks outperform unguided random mutation when the objective is to find under-constrained or mis-specified constraints  Ernstberger et al. (2024); Jiang et al. (2025).
@@ 2.3 LLMs and automated test/mutation generation
Large language models have been explored as code-generation and test-generation oracles across multiple domains. In compiler and test-generation settings, LLMs can be guided to produce high-quality test programs or mutation patterns when fed targeted prompts and feedback, enabling discovery of deep optimization and correctness bugs  Yang et al. (2024); Luo et al. (2021); Chen et al. (2021). Recent work combines LLMs with mutation testing and iterative refinement to generate tests that kill mutants and to produce sampling strategies tailored to observed failures  Hassan et al. (2024); Harman et al. (2025); Straubinger et al. (2025). Studies that evaluate LLMs specifically for ZK program generation report that off-the-shelf models often produce syntactically plausible but semantically incorrect ZK code, motivating augmentation strategies such as constraint sketching and guided retrieval  Xue et al. (2025); Li et al. (2025). These findings motivate zkCraft’s design choice to treat LLMs as deterministic template oracles rather than as formal solvers  Xue et al. (2025); Li et al. (2025).
@@ 2.4 Symbolic execution, SMT guidance, and hybrid strategies
Symbolic execution enhanced with solver-time prediction and synthesized solver strategies has been shown to improve scalability for complex programs, and these ideas inform hybrid pipelines that combine symbolic techniques with search or sampling  Luo et al. (2021); Chen et al. (2021); Trabish and Itzhaky (2026). For ZK circuit mutation and verification, solver-guided fallback channels that ground small template families and use finite-field SMT/SAT queries remain practical when pure ZK-native IOPs are infeasible; this hybrid approach balances completeness and engineering constraints  Bailey and Miller (2024); Almeida et al. (2021). Machine-checked proofs and formally verified protocol implementations provide strong guarantees and serve as an orthogonal means to reduce the trusted code base for ZK constructions  Almeida et al. (2021); Bailey and Miller (2024).
@@ 2.5 ZK-native proof techniques, commitments, and succinct verification
Recent advances in polynomial commitments, IOPs and succinct proof systems enable encoding higher-level search oracles as compact ZK statements. Systems that adopt ZK-native encodings show how commitment-and-IOP primitives can be repurposed to certify existential properties about edits and witnesses, thereby converting existential search into a single verifiable artifact  Mouris and Tsoutsos (2021); Bellés-Muñoz et al. (2022); Angel et al. (2022). GPU-accelerated proof systems and hardware-algorithm co-design work illustrate routes to practical prover performance that scale to larger circuits  Ma et al. (2023); Butt et al. (2024); Samardzic et al. (2024). Benchmarks and tooling that provide comparative performance baselines for SNARK stacks are useful for choosing commit/IOP backends when building ZK-native workflows  Ernstberger et al. (2024); Liu et al. (2025).
@@ 2.6 Circuit languages, compilers and verification frameworks
Domain-specific languages for ZK, their compiler frontends, and associated verification toolchains form a rich ecosystem. Efforts to improve circuit description languages and make them amenable to static verification encourage safer ZK application development  Bellés-Muñoz et al. (2022); Ozdemir et al. (2022). Compiler-focused fuzzing and white-box analysis have uncovered deep logic bugs in real-world compilers, motivating tighter integration of source-level diagnostics with pipeline-level checks  Yang et al. (2024); Xiao et al. (2025). Tools that certify circuits with refinement types or other higher-level specifications reduce a class of errors by construction and complement testing-based approaches  Liu et al. (2024); Jiang et al. (2025).
@@ 2.7 Applications, surveys, and domain-specific ZK use cases
Zero-knowledge proofs are being applied widely, from verifiable decentralized machine learning to privacy-preserving credentialing and domain-specific verification tasks  Xing et al. (2025); Berrios Moya et al. (2025); Gangurde et al. (2025). Surveys and systematizations contextualize verifiability challenges and catalog practical design patterns and open problems for ZKP-based applications  Lavin et al. (2024); Liang et al. (2025). Domain work that integrates anomaly detection, sensing, or constrained hardware contexts highlights practical verification needs in applied settings  Salam et al. (2024); Park et al. (2024); Akgul et al. (2025).
@@ 2.8 Mutation tooling, scientific debugging and cross-domain testing
Mutation testing and scientific debugging have been adapted to multiple domains including quantum circuits and hardware verification; these adaptations emphasize careful mutant design and oracle selection to obtain meaningful faults  Gil et al. (2024); Osama et al. (2025). Iterative, model-guided debugging pipelines that combine LLMs with mutation testing demonstrate that principled human-in-the-loop strategies and automated refinement can improve fault detection while keeping developer review effort manageable  Hassan et al. (2024); Harman et al. (2025); Straubinger et al. (2025). These cross-domain lessons guided the design of zkCraft’s mutation templates and deterministic post-processing to keep generated edits syntactically valid and semantically focused  Xue et al. (2025); Hassan et al. (2024).
@@ 2.9 Summary of positioning
In summary, prior work supplies three families of building blocks for zkCraft. The first family covers ZK pipeline correctness, compilers, and static/dynamic detectors that expose semantic vulnerabilities  Wen et al. (2024); Xiao et al. (2025); Hochrainer et al. (2025a); Li et al. (2024). The second family comprises mutation, fuzzing, and LLM-guided test generation techniques that provide high-yield input and edit proposals  Yang et al. (2024); Hassan et al. (2024); Harman et al. (2025); Straubinger et al. (2025). The third family contains ZK-native proof primitives, commitment/IOP engineering, and performance tooling that enable replacing many solver calls with compact cryptographic artifacts  Mouris and Tsoutsos (2021); Angel et al. (2022); Ma et al. (2023); Liu et al. (2025). zkCraft integrates elements from all three families by combining algebraic localization, a Row-Vortex commitment with a Violation IOP, and deterministic LLM-driven mutation templates to produce proof-bearing counterexamples that are both practical and auditable  Takahashi et al. (2025); Jiang et al. (2025); Li et al. (2025). guarantees, advancing the state-of-the-art in ZKP security.
@@ 3 Methodology
This section formalizes the zkCraft methodology and its ZK-native extension. We present a deterministic, R1CS-aware pipeline that replaces randomized program mutation with a small-scope, algebraic slicing procedure. The novel extension replaces the external SMT/SAT oracle with a succinct zero-knowledge protocol. Under this design the existence of a small edit set and the concrete edits themselves are certified by a single proof $\pi$; proof verification both confirms the vulnerability and yields the concrete counterexample trace without a separate TCCT Takahashi et al. (2025) invocation.
@@ 3.13 Implementation
Constraint #Bench #Bug Circomspect Blum et al. (2019) ZKAP Wen et al. (2024) Picus Pailoor et al. (2023) ConsCS Jiang et al. (2025) ZKFUZZ (Max / Min) Takahashi et al. (2025) ZKFUZZ++ Takahashi et al. (2025) zkCraft
@@
Type ZKFUZZ Takahashi et al. (2025) zkCraft
@@
CIVER Isabel et al. (2024) No
CODA Liu et al. (2024) No
Constraint Checker Fan et al. (2024) No
Circomspect Blum et al. (2019) $\circ$ $\circ$ Yes
Picus Pailoor et al. (2023) $\circ$ No
ConsCS Jiang et al. (2025) $\circ$ No
AC4 Chen et al. (2024) $\circ$ $\circ$ No
ZKAP Wen et al. (2024) $\circ$ $\circ$ Yes
ZKFUZZ Takahashi et al. (2025) No
ZKFUZZ++ Takahashi et al. (2025) No
@@ 4.1 Benchmarks and experimental setup
Baseline tools in the comparison are Circomspect Blum et al. (2019), ZKAP Wen et al. (2024), Picus Pailoor et al. (2023) (with Z3 and CVC5 backends), ZKFUZZ Takahashi et al. (2025) and
ConsCS Jiang et al. (2025). Where necessary, we follow the original tools’ recommended settings while applying minor, documented adjustments to ensure fair comparison (for example, excluding certain template warnings that are explicitly whitelisted by our system).
@@ 4.11 Additional Analysis and Discussion
We compare zkCraft with ZKFUZZ  Takahashi et al. (2025) and ZKAP  Wen et al. (2024) under an identical two-hour budget, hardware, and seeds. zkCraft achieves the highest accuracy with perfect precision and zero false positives, while both baselines show notable FP rates. Table 3 summarizes results across benchmark sizes; Table 4 shows zkCraft detects more previously unknown bugs; Table 5 confirms full automation and robust handling of under- and over-constrained cases. Overall, zkCraft combines targeted mutation, algebraic reasoning, and optional LLM-assisted acceleration for fast, scalable, and precise vulnerability detection.

### v2

@@ 1 Introduction
Zero-knowledge circuits provide succinct, verifiable proofs while exposing minimal information about private inputs, and they form the foundation of many privacy and scalability systems. Writing correct circuits is difficult because the witness generator that computes private values must align precisely with the constraint system used for proof verification. When this alignment fails, two broad classes of faults appear. Under-constrained faults arise when the constraint system admits execution traces that no valid witness can produce. Over-constrained faults arise when a valid witness-producing execution is rejected by the constraints. Both fault classes have important security and operational consequences in deployed systems and have driven significant recent research efforts  [41, 55, 14, 56, 44, 34].
Prior work addresses these problems from multiple perspectives. Static and pattern-based detectors reveal common syntactic mistakes but often miss semantic discrepancies that manifest only during witness generation or under specific inputs  [13, 53, 49]. Formal verification and SMT-based analyses can provide strong guarantees but typically do not scale to large circuits, and they struggle with the combined search space of program edits and input assignments  [54, 47, 24]. Dynamic approaches based on fuzzing and metamorphic testing improve practical bug-finding coverage, yet they frequently depend on many expensive solver calls or unguided mutations that produce numerous low-value candidates  [41, 55, 56]. Recent work on verifier-side encodings, interactive oracle proofs, and efficient polynomial commitments offers promising primitives to compress or certify search outcomes  [44, 64, 65]. In parallel, advances in model-guided mutation and LLM-driven fuzzing have shown that language models can generate semantically rich test patterns and seed mutations for security testing  [37, 62, 33, 63, 4, 29].
Despite these developments, important gaps remain. Most algebraic and IOP-style techniques have not been integrated with practical program-level mutation strategies, so solver costs remain a bottleneck for end-to-end bug discovery. Conversely, LLM-guided mutation methods improve suggestion quality but they are typically detached from proof-bearing search and therefore can produce proposals that are expensive to validate or that lead to many false positives  [50, 29, 62]. Compilers and zkVM toolchains introduce additional complexity related to optimization and backend behavior, which further complicates end-to-end detection and increases the cost of exhaustive solver-based validation  [18, 59, 66, 57].
@@ 2.1 Correctness and verification of ZK pipelines and compilers
The correctness of zero-knowledge compilation and circuit-processing pipelines has attracted growing attention because compiler or pipeline bugs can yield accepted yet invalid proofs. Static and dynamic analyses specifically targeting ZK circuit representations and their compilers have been proposed to reveal semantic and logic flaws early in the toolchain. Prior work presents static detectors over circuit dependence graphs and evaluates them on large Circom codebases, demonstrating that semantic checks can recover subtle vulnerabilities in real projects  [55]. Metamorphic and mutation-driven approaches have been adapted to the ZK compiler setting to detect incorrect compilation behaviors by generating semantically equivalent variants and checking for inconsistent outputs  [56, 28]. Systematic fuzzing of end-to-end ZK pipelines has also been shown to surface logic bugs that elude conventional testing, illustrating the importance of domain-aware test oracles for pipelines that span DSL, compiler, and proof-generation layers  [22, 10].
@@ 2.2 Fuzzing, metamorphic testing, and oracles for ZK systems
Fuzzing and metamorphic testing have been successfully adapted to find pipeline-level and finalization bugs in ZK deployments, including rollup finalization failures and pipeline logic flaws; these techniques emphasize careful oracle design and input-generation strategies to trigger diverse behaviors  [28, 43]. Tools that combine metamorphic relations with fault-injection or targeted mutation expose both soundness and completeness weaknesses in zkVMs and related infrastructures  [23, 28]. Work that specifically targets the oracle problem in ZK fuzzing highlights the need for proofs or semantic checks that distinguish genuine counterexamples from benign variations  [10, 22]. Complementary studies show that bounded, structured mutation templates and algebraic checks outperform unguided random mutation when the objective is to find under-constrained or mis-specified constraints  [15, 25].
@@ 2.3 LLMs and automated test/mutation generation
Large language models have been explored as code-generation and test-generation oracles across multiple domains. In compiler and test-generation settings, LLMs can be guided to produce high-quality test programs or mutation patterns when fed targeted prompts and feedback, enabling discovery of deep optimization and correctness bugs  [61, 35, 12]. Recent work combines LLMs with mutation testing and iterative refinement to generate tests that kill mutants and to produce sampling strategies tailored to observed failures  [21, 20, 48]. Studies that evaluate LLMs specifically for ZK program generation report that off-the-shelf models often produce syntactically plausible but semantically incorrect ZK code, motivating augmentation strategies such as constraint sketching and guided retrieval  [60, 27]. These findings motivate zkCraft’s design choice to treat LLMs as deterministic template oracles rather than as formal solvers  [60, 27].
@@ 2.4 Symbolic execution, SMT guidance, and hybrid strategies
Symbolic execution enhanced with solver-time prediction and synthesized solver strategies has been shown to improve scalability for complex programs, and these ideas inform hybrid pipelines that combine symbolic techniques with search or sampling  [35, 12, 52]. For ZK circuit mutation and verification, solver-guided fallback channels that ground small template families and use finite-field SMT/SAT queries remain practical when pure ZK-native IOPs are infeasible; this hybrid approach balances completeness and engineering constraints  [5, 2]. Machine-checked proofs and formally verified protocol implementations provide strong guarantees and serve as an orthogonal means to reduce the trusted code base for ZK constructions  [2, 5].
@@ 2.5 ZK-native proof techniques, commitments, and succinct verification
Recent advances in polynomial commitments, IOPs and succinct proof systems enable encoding higher-level search oracles as compact ZK statements. Systems that adopt ZK-native encodings show how commitment-and-IOP primitives can be repurposed to certify existential properties about edits and witnesses, thereby converting existential search into a single verifiable artifact  [38, 6, 3]. GPU-accelerated proof systems and hardware-algorithm co-design work illustrate routes to practical prover performance that scale to larger circuits  [36, 9, 46]. Benchmarks and tooling that provide comparative performance baselines for SNARK stacks are useful for choosing commit/IOP backends when building ZK-native workflows  [15, 32].
@@ 2.6 Circuit languages, compilers and verification frameworks
Domain-specific languages for ZK, their compiler frontends, and associated verification toolchains form a rich ecosystem. Efforts to improve circuit description languages and make them amenable to static verification encourage safer ZK application development  [6, 40]. Compiler-focused fuzzing and white-box analysis have uncovered deep logic bugs in real-world compilers, motivating tighter integration of source-level diagnostics with pipeline-level checks  [61, 56]. Tools that certify circuits with refinement types or other higher-level specifications reduce a class of errors by construction and complement testing-based approaches  [31, 25].
@@ 2.7 Applications, surveys, and domain-specific ZK use cases
Zero-knowledge proofs are being applied widely, from verifiable decentralized machine learning to privacy-preserving credentialing and domain-specific verification tasks  [58, 7, 17]. Surveys and systematizations contextualize verifiability challenges and catalog practical design patterns and open problems for ZKP-based applications  [26, 30]. Domain work that integrates anomaly detection, sensing, or constrained hardware contexts highlights practical verification needs in applied settings  [45, 42, 1].
@@ 2.8 Mutation tooling, scientific debugging and cross-domain testing
Mutation testing and scientific debugging have been adapted to multiple domains including quantum circuits and hardware verification; these adaptations emphasize careful mutant design and oracle selection to obtain meaningful faults  [19, 39]. Iterative, model-guided debugging pipelines that combine LLMs with mutation testing demonstrate that principled human-in-the-loop strategies and automated refinement can improve fault detection while keeping developer review effort manageable  [21, 20, 48]. These cross-domain lessons guided the design of zkCraft’s mutation templates and deterministic post-processing to keep generated edits syntactically valid and semantically focused  [60, 21].
@@ 2.9 Summary of positioning
In summary, prior work supplies three families of building blocks for zkCraft. The first family covers ZK pipeline correctness, compilers, and static/dynamic detectors that expose semantic vulnerabilities  [55, 56, 22, 28]. The second family comprises mutation, fuzzing, and LLM-guided test generation techniques that provide high-yield input and edit proposals  [61, 21, 20, 48]. The third family contains ZK-native proof primitives, commitment/IOP engineering, and performance tooling that enable replacing many solver calls with compact cryptographic artifacts  [38, 3, 36, 32]. zkCraft integrates elements from all three families by combining algebraic localization, a Row-Vortex commitment with a Violation IOP, and deterministic LLM-driven mutation templates to produce proof-bearing counterexamples that are both practical and auditable  [51, 25, 27]. guarantees, advancing the state-of-the-art in ZKP security.
@@ 3 Methodology
This section formalizes the zkCraft methodology and its ZK-native extension. We present a deterministic, R1CS-aware pipeline that replaces randomized program mutation with a small-scope, algebraic slicing procedure. The novel extension replaces the external SMT/SAT oracle with a succinct zero-knowledge protocol. Under this design the existence of a small edit set and the concrete edits themselves are certified by a single proof $\pi$; proof verification both confirms the vulnerability and yields the concrete counterexample trace without a separate TCCT [51] invocation.
@@ 3.13 Implementation
Constraint #Bench #Bug Circomspect [8] ZKAP [55] Picus [41] ConsCS [25] ZKFUZZ (Max / Min) [51] ZKFUZZ++ [51] zkCraft
@@
Type ZKFUZZ [51] zkCraft
@@
CIVER [24] No
CODA [31] No
Constraint Checker [16] No
Circomspect [8] $\circ$ $\circ$ Yes
Picus [41] $\circ$ No
ConsCS [25] $\circ$ No
AC4 [11] $\circ$ $\circ$ No
ZKAP [55] $\circ$ $\circ$ Yes
ZKFUZZ [51] No
ZKFUZZ++ [51] No
@@ 4.1 Benchmarks and experimental setup
Baseline tools in the comparison are Circomspect [8], ZKAP [55], Picus [41] (with Z3 and CVC5 backends), ZKFUZZ [51] and
ConsCS [25]. Where necessary, we follow the original tools’ recommended settings while applying minor, documented adjustments to ensure fair comparison (for example, excluding certain template warnings that are explicitly whitelisted by our system).
@@ 4.11 Additional Analysis and Discussion
We compare zkCraft with ZKFUZZ  [51] and ZKAP  [55] under an identical two-hour budget, hardware, and seeds. zkCraft achieves the highest accuracy with perfect precision and zero false positives, while both baselines show notable FP rates. Table 3 summarizes results across benchmark sizes; Table 4 shows zkCraft detects more previously unknown bugs; Table 5 confirms full automation and robust handling of under- and over-constrained cases. Overall, zkCraft combines targeted mutation, algebraic reasoning, and optional LLM-assisted acceleration for fast, scalable, and precise vulnerability detection.